OpenStudy (anonymous):

Write y = 2/3x + 7 in standard form using integers.

pooja195 (pooja195):

The standard form looks like: Ax + By = C

pooja195 (pooja195):

So, a good first step would be to get your initial equation in a form that doesn't have any fractions... y = –2/3x + 7 Try multiplying everything by 3.... what is your new equation then?

pooja195 (pooja195):

When I solve it, I get: y = –2/3x + 7 3y = -2x + 21 (Multiply through by 3) 2x + 31 = 21 (move 2x to the left side)

pooja195 (pooja195):

2x + 3y = 21

pooja195 (pooja195):

np :)

pooja195 (pooja195):

what problem dont you get post it here

pooja195 (pooja195):

@debbie17241

pooja195 (pooja195):

?????

OpenStudy (anonymous):

im here